# DReSS: Data-driven Regularized Structured Streamlining for Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have achieved significant progress across various domains, but their increasing scale leads to high computational and memory costs. Recent studies show that LLMs exhibit sparsity, which can be exploited for pruning. However, existing pruning methods typically follow a prune-then-finetune paradigm. Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, necessitating expensive fine-tuning to recover performance. To address this, we propose a new paradigm: first apply regularization, then prune, and finally fine-tune. Based on this paradigm, we introduce DReSS, a simple and effective Data-driven Regularized Structured Streamlining method for LLMs. By leveraging a small amount of data to regularize the components before pruning, DReSS transfers the important information to the remaining parts of the model in advance. Compared to direct pruning, this can reduce the information loss caused by parameter removal, thereby enhancing its language modeling capabilities. Experimental results demonstrate DReSS achieves comparable performance to conventional methods even without fine-tuning, drastically alleviating computational costs. Moreover, DReSS significantly outperforms existing pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have achieved significant advancements across a wide range of tasks and domains, demonstrating their robust capabilities (Zhang et al., 2022; Achiam et al., 2023; Touvron et al., 2023; Wu et al., 2024). However, as the model size increases, the growing number of parameters leads to significant computational and memory requirements, which significantly hinder the practical deployment of LLMs. Consequently, it is urgent to develop methods that can reduce model size while maintaining performance.
@@
To address these challenges, several methods have been proposed, including pruning (Frantar & Alistarh, 2023; Sun et al., 2023; Ma et al., 2023; An et al., 2024), quantization (Frantar et al., 2022; Xiao et al., 2023), knowledge distillation (Shridhar et al., 2022; Hsieh et al., 2023), and low-rank decomposition (Saha et al., 2023). In this work, we mainly focus on pruning—an efficient and highly generalizable approach that can be seamlessly integrated with other model compression strategies. Pruning techniques are generally classified into two primary categories: unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and structured pruning (Ma et al., 2023; An et al., 2024). Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) for specific downstream tasks without relying on specialized hardware (Zhu et al., 2024). Moreover, the model obtained through structured pruning typically achieves much faster inference speed due to the regular data patterns.
@@
Despite these advancements, existing structured pruning methods still have some limitations. They all follow the paradigm of first selecting channels or layers to prune based on a designed metric, and then performing RFT (Chavan et al., 2024). However, they neglect the fact that important information can also exist in the pruned parts (Dettmers et al., 2022; Xiao et al., 2023; Yin et al., 2023), and directly removing them leads to an irreversible decline in model performance. Thus, achieving satisfactory performance with pruned models often requires extensive data for RFT (Ma et al., 2023), increasing computational costs. Additionally, high pruning ratios frequently cause performance collapse, limiting their effectiveness in reducing latency and improving throughput.
@@
To tackle these limitations, in this paper, we propose a new pruning paradigm that first applies regularization to the components to be pruned, then removes those parameters, and finally performs RFT. To the best of our knowledge, we are the first to propose this paradigm. The regularization process explicitly transfers important information from the parts to be pruned to the remaining parts, thereby reducing the information loss caused by direct parameter removal. Based on this paradigm, we introduce DReSS, a Data-driven Regularized Structured Streamlining method for LLMs. The comparison of the proposed method with existing channel-wise pruning approaches is shown in Figure 1. Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrix, and (3) prune the specified channels, finally (4) do RFT. First, we randomly select a small subset of data from widely used benchmark datasets. Since the amount of selected data is small, the overhead in the regularization and RFT process can be significantly reduced. Second, we determine the channels to be pruned based on the pruning ratio, and apply regularization to these channels using the selected data. This redistributes important information to the unpruned components, thereby significantly mitigating performance degradation caused by direct pruning. As a result, the model can maintain robust performance even at high pruning ratios, facilitating a substantial speedup in inference. Third, we prune the parameters that have been regularized. Finally, we perform RFT on the pruned model using a small subset of samples selected in the first step. Extensive experiments show that DReSS significantly outperforms existing pruning methods in perplexity and accuracy across various pruning ratios, even without RFT, while also greatly enhancing inference speed. The main contributions of our work are summarized as follows:
@@
  • Propose A New Paradigm: By sequentially applying regularization, pruning, and RFT, DReSS minimizes information loss caused by direct parameter removal, thereby improving the model’s language modeling capabilities and reducing the computational cost of RFT.
  • High Performance: DReSS surpasses existing structured pruning methods in both generation and zero-shot tasks. The pruned model achieves a substantial increase in throughput and a significant reduction in latency compared to the dense model.
@@
  • Low Overhead: By leveraging only a small amount of data for regularization and optional RFT, DReSS incurs minimal overhead.
@@
Figure 1: A comparison between existing channel-wise structured pruning methods and DReSS. The deeper blue square represents greater performance impact; the red exclamation mark indicates the important information. The taller orange-yellow cylinder represents larger data volume. L($\cdot$) represents the regularization, and S represents pruning. Above: Existing methods first select channels based on importance, followed by pruning and RFT on the pruned model. Below: DReSS first regularizes channels to transfer important information, prunes the channels, then do RFT with a little amount of data.
@@ 2.1 Pruning Methods
The elimination of redundant weights, referred to as pruning, has been an effective approach for reducing the complexity of deep neural networks over the past several decades (LeCun et al., 1989; Hassibi et al., 1993; Han et al., 2015; Wen et al., 2016; Louizos et al., 2017). Pruning methods can be broadly categorized into two types: unstructured pruning (Kurtic et al., 2022; Zhang et al., 2024; Xu et al., 2024) and structured pruning (Xia et al., 2023b; Gao et al., 2024b; Yang et al., 2024).
Unstructured pruning methods remove individual weights based on their importance. For instance, Magnitude (Han et al., 2015; Zhou et al., 2021) assumes larger weights are more important, Wanda (Sun et al., 2023) uses the product of layer weights and activations as the metric, and SparseGPT (Frantar & Alistarh, 2023) relies on the weight squared divided by the inverse Hessian matrix. The main advantage is flexibility, enabling the removal of unimportant weights while retaining critical ones, allowing high sparsity in LLMs. However, its limitations include the requirement for specialized hardware to accelerate inference (Xia et al., 2023a), the necessity for high sparsity to achieve substantial acceleration (Wang, 2020), and the inability to perform RFT on downstream tasks.
Structured pruning methods, including channel-wise pruning (An et al., 2024) and layer-wise pruning (Men et al., 2024), can achieve significant inference speedup with a low sparsity. Channel-wise pruning methods typically design a metric to evaluate the importance of channels in the parameter matrix of LLMs and then prune the less important channels. For instance, SliceGPT (Ashkboos et al., 2024) adopts Principal Component Analysis (PCA), using the eigenvalues of the covariance matrix of activation values as a metric to preserve channels corresponding to larger eigenvalues. LLM Surgeon (van der Ouderaa et al., 2023) periodically updates model weights and structures, pruning more aggressively in the initial layers and less so in the intermediate layers. Layer-wise pruning methods, such as SLEB (Song et al., 2024), iteratively prune entire transformer layers by assessing their importance. However, these methods have a key limitation: even less important channels or layers may contain valuable information, and pruning them directly often leads to significant performance degradation. To address this, we investigate the possibility of ‘shifting’ important information from the pruned parts to the retained parts, which could substantially improve model performance.
@@ 2.2 Regularization
Regularization is widely employed in machine learning (Hoerl & Kennard, 1970; Poggio et al., 1987; Balestriero et al., 2022), such as feature selection (Tibshirani, 1996) and preventing model overfitting (Santos & Papa, 2022). The $\ell_{1}$-norm typically induces sparsity by driving certain coefficients to zero, while $\ell_{2}$-norm encourages smoother solutions (Boyd & Vandenberghe, 2004). Both of them can significantly alter the distribution pattern of the data (Han et al., 2015; Tao et al., 2023). Motivated by this observation, we propose a data-driven regularization approach. By applying a small amount of data, the regularization transfers important information from the pruned parameter space to the retained parts of the model, thereby enhancing the performance of the pruned model.
@@ 3 Methodology
In this section, we provide a detailed introduction to DReSS. The overall DReSS algorithm is summarized in Algorithm 1, our method consists of four main steps:
@@
  • Regularization: Applying regularization to the selected channels of weight matrices during forward and backward propagations.
  • Pruing: Prune the regularized portions.
  • RFT: Perform RFT on the pruned model using the subset of data selected in the first step.
@@
  Input: selected data $X$, number of layers in LLMs $l$, initial model $W$, number of regularization iterations $n$, norm type: flag.
  if flag==1 then
     use Proposition 3.2 to transform problem
@@
  $X_{1},X_{2}\leftarrow divide(X)$
  for $i=1$ to $n$ do
@@
     ~$\tilde{\mathcal{L}}_{\text{att}}$$\leftarrow$00, ~$\tilde{\mathcal{L}}_{\text{FFN}}$$\leftarrow$00, ~$\tilde{\mathcal{L}}_{\text{2}}$$\leftarrow$00
     for $j=1$ to $l$ do
        ~$\tilde{\mathcal{L}}_{\text{att}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{att}}$+$regularize(W_{Q}^{j},W_{K}^{j},W_{V}^{j},W_{o}^{j},X_{1})$
        ~$\tilde{\mathcal{L}}_{\text{FFN}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{FFN}}$+$regularize(W_{up}^{j},W_{down}^{j},X_{1})$
        ~$\tilde{\mathcal{L}}_{\text{2}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{2}}$+$regularize(w_{N}^{j},X_{1})$
     end for
[…]
     ~$\tilde{\mathcal{L}}_{\text{remain}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{1}}$+~$\tilde{\mathcal{L}}_{\text{2}}$+~$\tilde{\mathcal{L}}_{\text{3}}$
     $\mathcal{L}(W,X)\leftarrow forward(W,X_{1})$
     $\mathcal{L}_{\text{sum}}$$\leftarrow$$\mathcal{L}(W,X)$+~$\tilde{\mathcal{L}}_{\text{att}}$+~$\tilde{\mathcal{L}}_{\text{FFN}}$+~$\tilde{\mathcal{L}}_{\text{remain}}$
     update $W$ using backpropagation algorithm
  end for
  $W\leftarrow Prune(W,R)$
  $W\leftarrow Optional\_RFT(W,X_{2})$
3.1 Data Selection
The required data is selected from widely used calibration datasets, including Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020). For instance, approximately 1,000 samples are randomly selected from the WikiText-2 training dataset for pre-pruning regularization and post-pruning RFT. In Section 4.5, we analyze the impact of using different datasets.
3.2 Regularization
To provide a clearer understanding of this section, we first introduce the relevant notations. Let $p$ denote the pruning ratio, which indicates that $p\%$ of the model’s parameters need to be pruned. The model dimension of the LLMs is represented by $d$, and the number of layers is denoted by $l$. Additionally, let $W_{emb}\in\mathbb{R}^{b\times n\times d}$ represents the data feature map, and $W_{pos}\in\mathbb{R}^{b\times n\times d}$ denotes the positional feature map, where $b$ is the batch size and $n$ is the number of tokens. The model weights are represented by $W\in\mathbb{R}^{d_{1}\times d_{2}}$. Specifically, $W_{Q}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{K}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{V}^{i}\in\mathbb{R}^{d\times d_{1}}$, and $W_{o}^{i}\in\mathbb{R}^{d_{1}\times d}$ represent the parameter matrices in the $i$th layer Attention block. Similarly, $W_{up}^{i}\in\mathbb{R}^{d\times d_{2}}$ and $W_{down}^{i}\in\mathbb{R}^{d_{2}\times d}$ represent the parameter matrices in the $i$th layer FFN block. Additionally, $w_{N}^{i}\in\mathbb{R}^{d}$ denotes the LayerNorm vector in the $i$th layer, and $W_{lm}\in\mathbb{R}^{d\times b\times n}$ represents the language modeling matrix. Finally, $I\in\mathbb{R}^{d\times d}$ represents the identity matrix, let $R\in\mathbb{R}^{d\times d}$ represent a pseudo-index selection matrix, which is a diagonal matrix consisting of $(1-p\%)\times d$ zeros and $p\%\times d$ ones, where the ones correspond to the selected indices.
Figure 2: The pruning strategy of DReSS in a transformer layer.
Proposition 3.1.
(Dependency between the Attention Block and FFN Block. Proof in Appendix A.1). If regularization is applied to certain columns of $W_{down}^{i-1}$ in the $(i-1)$th layer FFN block, then regularization must also be applied to the corresponding rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th layer Attention block. The same dependency applies to $W_{o}^{i}$ and $W_{up}^{i-1}$.
We illustrate the process of applying regularization within a transformer layer (Vaswani, 2017), as shown in Figure 2. Following Proposition 3.1, in the Attention block, regularization is applied to the rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ corresponding to the one entries in the index matrix $R$, as well as to the columns of $W_{o}^{i}$ associated with these entries. Similarly, in the FFN block, regularization is enforced on the rows of $W_{up}^{i}$ and the columns of $W_{down}^{i}$ that correspond to the one entries in $R$. We also apply regularization to the remaining parts in LLMs. Specifically, for $W_{emb}$, regularization is enforced on the columns corresponding to the one indices in $R$, and the same procedure is applied to $W_{pos}$. For each $w_{N}^{i}$ within every transformer layer, regularization is applied to the elements corresponding to the one indices in $R$. Additionally, for the final $W_{lm}$, due to its dependency on $W_{down}^{l}$ in the FFN block of the last transformer layer, regularization is applied to the corresponding rows. Therefore, the overall loss function consists of the language modeling loss and the loss induced by regularization, which is defined below:
$\begin{split}\mathcal{L}_{\text{sum}}=\mathcal{L}(W,X)+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}+\tilde{\mathcal{L}}_{\text{remain}}\end{split}$ (1)
where $\mathcal{L}(W,X)$ denotes the language modeling loss and $X$ is the input data, and the remaining term represents the regularization loss. ~$\tilde{\mathcal{L}}_{i}^{\text{att}}$ denotes the $i$th layer attention block regularization loss, which is defined as:
~$\begin{split}\tilde{\mathcal{L}}_{i}^{\text{att}}=\lambda(\|RW_{Q}^{i}\|+\|RW_{K}^{i}\|+\|RW_{V}^{i}\|+\|W_{o}^{i}R\|)\end{split}$ (2)
~$\tilde{\mathcal{L}}_{i}^{\text{FFN}}$ denotes the $i$th layer FFN block regularization loss and the definition is as below:
~$\begin{split}\tilde{\mathcal{L}}_{\text{FFN}}^{i}=\lambda\|RW_{up}^{i}\|+\lambda\|W_{down}^{i}R\|\end{split}$ (3)
~$\tilde{\mathcal{L}}_{\text{remain}}$ denotes the remaining parts of the LLMs’ regularization loss as follows:
~$\begin{split}\tilde{\mathcal{L}}_{\text{remain}}=\lambda(\|W_{emb}R\|+\|W_{pos}R\|+\\ \sum_{i=1}^{l}\|Rw_{N}^{i}\|+\|RW_{lm}\|)\end{split}$ (4)
[…]
Proposition 3.2.
@@ 3.2 Regularization
(If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation. Proof in Appendix A.2). The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.

### v3

@@ Abstract
Large language models (LLMs) have achieved significant progress across various domains, but their increasing scale leads to high computational and memory costs. Recent studies show that LLMs exhibit sparsity, which can be exploited for pruning. Existing pruning methods typically follow a prune-then-finetune paradigm. Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, causing expensive fine-tuning to recover performance. To address this, we propose a new paradigm: first apply regularization, then prune, and finally fine-tune. Based on this paradigm, we propose DReSS, a simple and effective Data-driven Regularized Structured Streamlining method for LLMs. By using a small amount of data to regularize the components before pruning, DReSS transfers the important information to the remaining parts of the model in advance. Compared to direct pruning, this can reduce the information loss caused by parameter removal, thereby enhancing its language modeling capabilities. We evaluate our method on various LLMs, including Phi-2, OPT, LLaMA2, LLaMA3. Experimental results demonstrate DReSS even without recovery fine-tuning (RFT) achieves comparable performance to previous methods, drastically alleviating computational costs. Moreover, DReSS significantly outperforms existing powerful pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
@@ 1 Introduction
Large language models (LLMs) have achieved significant advancements across a wide range of tasks, demonstrating their robust capabilities (Zhang et al. 2022; Achiam et al. 2023; Touvron et al. 2023; Wu et al. 2024). However, as the model size increases, the growing number of parameters leads to significant computational and memory requirements, which significantly hinder the practical deployment of LLMs. Consequently, it is critical to develop methods that can reduce model size while maintaining performance.
@@
To address these challenges, several methods have been proposed, including pruning (Frantar and Alistarh 2023; Sun et al. 2024; An et al. 2024), quantization (Frantar et al. 2023; Xiao et al. 2023), knowledge distillation (Shridhar, Stolfo, and Sachan 2023; Hsieh et al. 2023), and low-rank decomposition (Saha, Srivastava, and Pilanci 2023). In this work, we mainly focus on pruning which is an efficient and highly generalizable approach that can be seamlessly integrated with other model compression strategies. Pruning techniques are generally classified into two primary categories: unstructured pruning (Frantar and Alistarh 2023; Sun et al. 2024) and structured pruning (Ma, Fang, and Wang 2023; An et al. 2024). Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) for specific downstream tasks without relying on specialized hardware (Zhu et al. 2024). The model obtained through structured pruning typically achieves much faster inference speed due to the regular data patterns.
@@
Despite these advancements, existing structured pruning methods still have limitations. They all follow the paradigm of first selecting channels or layers to prune based on a designed metric, and then performing RFT (Chavan et al. 2024). However, they neglect that important information can also exist in the pruned parts (Dettmers et al. 2022; Xiao et al. 2023; Yin et al. 2024), and directly removing them leads to an irreversible decline in model performance. Thus, the pruned models often require extensive data for RFT to recover performance (Ma, Fang, and Wang 2023). Additionally, high pruning ratios often cause performance collapse, limiting their effectiveness in acceleration.
@@
To tackle these limitations, we propose a new pruning paradigm that first applies regularization to the components to be pruned, then prunes those parameters, finally performs RFT. To the best of our knowledge, we are the first to propose this paradigm. The regularization process explicitly transfers important information from the parts to be pruned to the remaining parts, thereby reducing the information loss caused by direct parameter removal. Based on this paradigm, we introduce DReSS, a Data-driven Regularized Structured Streamlining method for LLMs. The comparison of the DReSS with existing channel-wise pruning approaches is shown in Figure 1. Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrices, (3) prune the regularized channels, finally (4) do RFT. First, we randomly select a small subset of data from widely used benchmark datasets. As the data size is small, the overhead of regularization and RFT remains minimal. Second, we determine the channels to be pruned based on the pruning ratio, and apply regularization ($\ell_{1}$-norm or $\ell_{2}$-norm) to these channels using the selected data. This redistributes important information to the unpruned parts, thereby significantly mitigating performance degradation caused by direct pruning. As a result, the model retains strong performance at high pruning ratios, enabling significant speedup. Third, we prune the regularized channels. Finally, we perform RFT on the pruned model using a small subset of samples selected in the first step. Extensive experiments show that DReSS significantly outperforms existing powerful pruning methods in perplexity and accuracy across various pruning ratios on different models and is able to provide considerable acceleration. The main contributions are summarized as follows:
@@
  • Propose A New Paradigm: By sequentially applying regularization ($\ell_{1}$-norm or $\ell_{2}$-norm), pruning, and RFT, DReSS minimizes information loss caused by direct parameter removal, thereby improving the model’s language modeling capabilities.
  • High Performance: DReSS surpasses competitive structured pruning methods in perplexity and accuracy, achieving notable improvements in throughput and reduced latency compared to the dense models.
@@
  • Low Overhead: By using only a small amount of data for regularization and optional RFT, DReSS incurs minimal overhead.
@@
Figure 1: A comparison between previous channel-wise pruning methods and DReSS. Deeper blue square represents greater performance impact, the taller cylinder represents larger data volume. Above: Previous methods first select channels based on importance, followed by pruning and then do RFT. Below: DReSS first regularizes channels to transfer important information, prunes the channels, then do RFT with a small amount of data.
@@ 2.1 Pruning Methods
Pruning redundant weights has been an effective way to reduce deep neural network complexity for decades. (LeCun, Denker, and Solla 1989; Hassibi, Stork, and Wolff 1993; Han et al. 2015). Pruning methods can be broadly categorized into two types: unstructured pruning (Kurtic et al. 2022; Zhang et al. 2024; Xu et al. 2024) and structured pruning (Xia et al. 2024; Gao et al. 2024b).
Unstructured pruning methods remove individual weights based on designed metrics. Magnitude (Han et al. 2015) removes smaller weights, Wanda (Sun et al. 2024) considers both weights and activations, and SparseGPT (Frantar and Alistarh 2023) uses Hessian matrix. They require specialized hardware to accelerate (Xia et al. 2023) and a high sparsity to achieve substantial acceleration (Wang 2020).
Structured pruning methods, including channel-wise pruning (An et al. 2024) and layer-wise pruning (Men et al. 2024). Channel-wise pruning methods remove less important channels of the parameter matrices. For instance, SliceGPT (Ashkboos et al. 2024) prune channels with smaller eigenvalues. LLM Surgeon (van der Ouderaa et al. 2024) periodically updates model weights and structures, pruning more aggressively in the initial layers. Layer-wise pruning methods, such as SLEB (Song et al. 2024), iteratively prune transformer layers based on their importance. These methods have a key limitation: even less important channels or layers may contain valuable information, and pruning them directly often leads to information loss. To address this, we investigate the way of ‘shifting’ important information from the pruned parts to the remaining parts, which could reduce information loss caused by pruning.
@@ 2.2 Regularization
Regularization is widely used in machine learning, such as feature selection (Tibshirani 1996) and preventing model overfitting (Santos and Papa 2022). The $\ell_{1}$-norm drives certain coefficients to zero, while $\ell_{2}$-norm encourages smoother solutions (Boyd and Vandenberghe 2004). Both of them can significantly alter the distribution pattern of the data (Han et al. 2015; Tao et al. 2023). Motivated by this, we propose a new pruning paradigm, in which the regularization process can transfer important information from the pruned parameter space to the remaining parts of the model, thus reducing the information loss caused by parameter removal.
@@ 3 Methodology
In this section, we provide a detailed introduction to DReSS. The overall DReSS algorithm is summarized in Figure 1 and Algorithm 1, our method consists of four main steps:
@@
  • Regularization: Apply regularization to the selected channels of weight matrices to transfer important information from the pruned parts to the remaining parts, thereby reducing information loss caused by pruning and mitigating the resulting performance degradation.
  • Pruing: Remove the channels that were regularized in the previous step.
  • RFT: Perform RFT on the pruned model to compensate for the performance loss caused by pruning.
3.1 Data Selection
For fairness, the data is selected from widely used datasets. For instance, approximately 1,000 samples are randomly selected from the WikiText-2 (Merity et al. 2016) training set for pre-pruning regularization and post-pruning RFT.
3.2 Regularization
For clarity, we first define the notations. $p$ is the pruning ratio, indicating $p\%$ of model parameters will be pruned. The model dimension is $d$, and the number of layers is $l$. Additionally, $\mathbf{W_{emb}}\in\mathbb{R}^{b\times n\times d}$ is the data feature map, and $\mathbf{W_{pos}}\in\mathbb{R}^{b\times n\times d}$ is the positional feature map, where $b$ is the batch size and $n$ is the number of tokens. The model weights are represented by $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$. Specifically, in the $i$th layer attention block, $\mathbf{W_{Q}^{i}}\in\mathbb{R}^{d\times d_{1}}$ is query matrix, $\mathbf{W_{K}^{i}}\in\mathbb{R}^{d\times d_{1}}$ is key matrix, $\mathbf{W_{V}^{i}}\in\mathbb{R}^{d\times d_{1}}$ is value matrix, and $\mathbf{W_{o}^{i}}\in\mathbb{R}^{d_{1}\times d}$ is output matrix. In the $i$th layer FFN block, $\mathbf{W_{up}^{i}}\in\mathbb{R}^{d\times d_{2}}$ is up projection matrix, and $\mathbf{W_{down}^{i}}\in\mathbb{R}^{d_{2}\times d}$ is down projection matrix. In the $i$th LayerNorm, $\mathbf{\alpha^{i}}\in\mathbb{R}^{d}$ is scaling vector and $\mathbf{\beta^{i}}\in\mathbb{R}^{d}$ is offset vector. $\mathbf{W_{lm}}\in\mathbb{R}^{d\times b\times n}$ denotes the language modeling matrix. $\mathbf{I}\in\mathbb{R}^{d\times d}$ is the identity matrix, let $\mathbf{R}\in\mathbb{R}^{d\times d}$ is a pseudo-index selection matrix, which is a diagonal matrix consisting of $(1-p\%)\times d$ zeros and $p\%\times d$ ones, where the ones correspond to the selected indices. $\mathbf{R}$ is identical across all layers.
@@ 3.2 Regularization
  Input: selected data $\mathbf{X}$, number of layers in LLMs $l$, initial model $\mathbf{W}$, norm type: flag.
  if flag==0 then
     continue// $\ell_{2}$-norm
  else if flag==1 then
     use Proposition 1 to transform problem // $\ell_{1}$-norm
@@
  $\mathbf{X_{1}},\mathbf{X_{2}}\leftarrow divide(\mathbf{X})$
  ~$\tilde{\mathcal{L}}_{\text{1}}\leftarrow regularize(\mathbf{W_{emb}},\mathbf{W_{pos}},\mathbf{X_{1}})$
  ~$\tilde{\mathcal{L}}_{\text{att}}\leftarrow 0$, ~$\tilde{\mathcal{L}}_{\text{FFN}}\leftarrow 0$, ~$\tilde{\mathcal{L}}_{\text{2}}\leftarrow 0$
  for $i=1$ to $l$ do
     ~$\tilde{\mathcal{L}}_{\text{att}}\leftarrow\tilde{\mathcal{L}}_{\text{att}}+regularize(Attention,\mathbf{X_{1}})$
     ~$\tilde{\mathcal{L}}_{\text{FFN}}\leftarrow\tilde{\mathcal{L}}_{\text{FFN}}+regularize(FFN,\mathbf{X_{1}})$
     ~$\tilde{\mathcal{L}}_{\text{2}}\leftarrow\tilde{\mathcal{L}}_{\text{2}}+regularize(LayerNorm,\mathbf{X_{1}})$
  end for
  ~$\tilde{\mathcal{L}}_{\text{3}}\leftarrow regularize(\mathbf{W_{lm}},\mathbf{X_{1}})$
  ~$\tilde{\mathcal{L}}_{\text{remain}}\leftarrow\tilde{\mathcal{L}}_{\text{1}}+\tilde{\mathcal{L}}_{\text{2}}+\tilde{\mathcal{L}}_{\text{3}}$
  $\mathcal{L}(\mathbf{W},\mathbf{X})\leftarrow forward(\mathbf{W},\mathbf{X_{1}})$
  $\mathcal{L}_{\text{sum}}\leftarrow\mathcal{L}(\mathbf{W},\mathbf{X})+\tilde{\mathcal{L}}_{\text{att}}+\tilde{\mathcal{L}}_{\text{FFN}}+\tilde{\mathcal{L}}_{\text{remain}}$
  update $\mathbf{W}$ using backpropagation algorithm
  $\mathbf{W}\leftarrow Prune(\mathbf{W},\mathbf{R})$
  $\mathbf{W}\leftarrow Optional\_RFT(\mathbf{W},\mathbf{X_{2}})$
Figure 2: The regularization and pruning process of DReSS in a transformer layer.
Proposition 1 (Dependency between attention and FFN block. Proof in Appendix A). If regularization is applied to certain columns of $\mathbf{W_{down}^{i-1}}$ in the $(i-1)$th layer FFN block, then regularization should also be applied to the corresponding rows of $\mathbf{W_{Q}^{i}}$, $\mathbf{W_{K}^{i}}$, and $\mathbf{W_{V}^{i}}$ in the $i$th layer attention block. The same dependency applies to $\mathbf{W_{o}^{i}}$ and $\mathbf{W_{up}^{i-1}}$.
Figure 2 illustrates the process of applying regularization in a transformer layer. Following Proposition 1, in the attention block, regularization is applied to the rows of $\mathbf{W_{Q}^{i}}$, $\mathbf{W_{K}^{i}}$, and $\mathbf{W_{V}^{i}}$ corresponding to the one entries in the index matrix $\mathbf{R}$ and the columns of $\mathbf{W_{o}^{i}}$ associated with these entries. In the FFN block, regularization is applied on the rows of $\mathbf{W_{up}^{i}}$ and the columns of $\mathbf{W_{down}^{i}}$ that correspond to the one entries in $\mathbf{R}$. We also apply regularization to the remaining parts in LLMs. Specifically, for $\mathbf{W_{emb}}$, regularization is enforced on the columns corresponding to the one indices in $\mathbf{R}$, and the same for $\mathbf{W_{pos}}$. For each $\mathbf{\alpha^{i}}$ and $\mathbf{\beta^{i}}$ within every transformer layer, regularization is applied to the elements corresponding to the one indices in $\mathbf{R}$. Additionally, for the final $\mathbf{W_{lm}}$, due to its dependency on $\mathbf{W_{down}^{l}}$ in the FFN block of the last transformer layer, regularization is applied to the corresponding rows. Therefore, the overall loss function defined in Equation 1 consists of the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$ and the regularization loss. $\mathbf{X}$ is the input data. Regularization loss has three parts: Attention loss: $\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}$, FFN loss: $\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}$, remain loss: ~$\tilde{\mathcal{L}}_{\text{remain}}$.
$\begin{split}\mathcal{L}_{\text{sum}}=\mathcal{L}(\mathbf{W},\mathbf{X})+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}+\tilde{\mathcal{L}}_{\text{remain}}\end{split}$ (1)
~$\tilde{\mathcal{L}}_{\text{att}}^{i}$ denotes the $i$th layer attention block regularization loss:
~$\begin{split}\tilde{\mathcal{L}}_{\text{att}}^{i}=\lambda(\|\mathbf{RW_{Q}^{i}}\|+\|\mathbf{RW_{K}^{i}}\|+\|\mathbf{RW_{V}^{i}}\|+\|\mathbf{W_{o}^{i}R}\|)\end{split}$ (2)
~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ is the $i$th layer FFN block regularization loss:
~$\begin{split}\tilde{\mathcal{L}}_{\text{FFN}}^{i}=\lambda\|\mathbf{RW_{up}^{i}}\|+\lambda\|\mathbf{W_{down}^{i}R}\|\end{split}$ (3)
~$\tilde{\mathcal{L}}_{\text{remain}}$ denotes the regularization loss of the remaining parts of the LLM:
~$\begin{split}\tilde{\mathcal{L}}_{\text{remain}}=\lambda(\|\mathbf{W_{emb}R}\|+\|\mathbf{W_{pos}R}\|+\\ \sum_{i=1}^{l}\|\mathbf{R\alpha^{i}}\|+\sum_{i=1}^{l}\|\mathbf{R\beta^{i}}\|+\|\mathbf{RW_{lm}}\|)\end{split}$ (4)
@@
Proposition 2 (If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation. Proof in Appendix B). The following two problems are equivalent, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.